A particle of mass 0.1kg is executing SHM of amplitude 1m and period 0.2 sec what max. Value of force acting on it? If oscillation is produced by spring what is the force constant of spring

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Maximum acceleration
a(max) = w² *a = 4\pi² A/ T² = 4 *(3.14)² * 1/(0.2)²
 
F(max) = m * a(max) = 0.1 * 4 * (3.14)²*1/(0.2)²
 = 98.6 N
 
Force constant k = Fmax/A = 98.6/1
 = 98.6 N/m
 
 
Regards
Arun (askIITians forum expert)